OpenStudy (anonymous):

from (2,1) to (5,2) is three units how many units is it from (5,1) to (5,-3)?

myininaya (myininaya):

d=sqrt((2-5)^2+(1-2)^2)=sqrt(9+1)=sqrt(10) ? are we talking about a length of line segment whose endpoints are (2,1) and (5,2)?

OpenStudy (anonymous):

i dont know thats what the question says